# Exploring the Effects of User-Agent and User-Designer Similarity in Virtual Human Design to Promote Mental Health Intentions for College Students

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TAPjournalvolume: Xjournalnumber: Xarticle: Xccs: Computing methodologies Intelligent agents
@@ 3.1. Participants
Participants’ reported ages were: 18-20 years of age (n=339), 21-25 years of age (n=119), 26-30 years of age (n=8), and over 30 years of age (n=13). Two participants (n=2) did not disclose their age. Participants also reported the number of languages they were familiar with–monolingual (n=197) or multilingual (n=284), and their place of origin–country and city. Participants reported being from 34 different countries: from the United States (n=355) or other places (n=126)–see Appendix A, Table 9. Participants also reported being domestic (n=402) or international (n=79) students.
@@ 3.3. Data analysis
Moreover, the study examined how participants selected virtual humans’ appearances and voices when prompted to choose options that would encourage ”someone similar to them” to engage in mental wellness conversations. Top preferences on appearances and voices were analyzed using descriptive statistics, cross-referenced with participants’ demographics. We used the Pearson correlation coefficient to compare participants’ self-reported ages with the perceived ages of the agents’ appearances they selected. Additionally, we calculated the Jaccard similarity coefficient to evaluate the relationship between participants’ self-reported ethnic groups and the ethnic groups they perceived the agents’ appearances they selected represented.
@@ 3.4. Findings: Agent-designer matching of demographic characteristics
Of the four characteristics matched between participants and virtual humans they co-designed, age was the only demographic characteristic for which agent-participant (i.e., agent-designer) matching was not the preferred choice among most participants when designing a virtual human to provide mental health assistance to someone similar to them. Three participants (n=3) did not report a counselor’s age. Among the remaining participants who self-reported their age (n=476), 66.9% of participants preferred an older virtual human (more than five years older), 0.4% preferred a younger virtual human (less than five years younger), and 32.8% designed a virtual human as old as they reported to be (having a window of ±5 years of age for similarity): virtual humans’ average age was of 28.6 years of age, reporting a maximum of 60 years of age. The Pearson correlation coefficient revealed a significant association between participants’ self-reported ages and the virtual agents’ ($r=0.257,n=476,p<0.01$), suggesting a positive relationship between them. This correlation suggests that as the participant’s age increases, the agent’s age also tends to increase, which supports the existence of user-agent age similarity when participants design agents to support individuals similar to them in mental well-being.
@@
From our sample population (n=481), 61.1% of participants decided on ethnic matching and 76.9% on matching their country of origin: 24.1% indicated the same city of origin. The sample population had male (n=281), female (n=182), and participants from other gender groups (n=14)–e.g., non-binary, transgender, etc. From the general population, 53.0% of them preferred a female agent, 36.0% a male one, and 9.8% did so for an agent with another gender identity. See Appendix A: Table 10 shows preferences on gender concerning participants’ ethnic groups.
Looking at gender similarity, 72.8% of the general population decided on gender matching. Female participants (n=182) highly preferred a female agent: 90.1% of the female population. Similarly, 78.6% of participants who identified with other gender identities (n=14–e.g., non-binary, gender-queer, transgender) preferred an agent who gender-matched them. Male participants (n=281) also leaned toward a gender-matching agent, although not as pronounced as participants from other gender groups (61.9% of males). Therefore, gender similarity is critical in the design of virtual humans to support non-Male minorities in mental health.
@@ 3.6.1. Perceptions of age and ethnicity of virtual human male appearances
Of the seven ethnic groups considered, two of them did not tend toward ethnic similarity concerning male appearances. Among Latin American participants (n=100), only 33.0% of them selected a male appearance they perceived as Latin American. Likewise, only 21.4% of Middle Eastern/North African participants (n=14) leaned toward ethnically similar male appearances. As 93.0% of the Latin American and 92.3% of the Middle Eastern/North African population were domestic, further research is needed to understand if this phenomenon is due to limitations on the catalog of appearances or familiarity aspects regarding the under-representation of these ethnic groups in counseling scenarios in the U.S.A.–counselors in the U.S.A. are predominantly White (non-Latin American) (USA, 2017).
@@ 3.9. Findings: Preferences on encouraging virtual human male voices
Similar to the findings reported by Feijóo-García et al. (Feijóo-García et al., 2023), there were voice preferences among Indian participants (n=57) based on user-agent matching regarding the place of origin. It was observed in participants’ responses when selecting the four most encouraging male voices: 63.2% of Indian participants (n=57) selected the male Indian voice (IN_M). However, this effect was not observed in their top one preference nor among participants from other places of origin.
@@ 3.10. Findings: Preferences on encouraging virtual human female voices
Similar to observations on participants’ choices of male voices (Section 3.9), Indian participants (n=57) leaned toward female voices that matched their place of origin when selecting the four most encouraging female voices: 80.7% of Indians (n=57) selected the female Indian voice (IN_F). Similarly, participants from the United States (n=354) generally included the American voice (US_F) in their four most encouraging female voices (56.2% of responses). Nevertheless, as with male voice selections, this effect was not observed in participants’ final top preferences for female voices.
@@
Accent similarity expectations were not observed among international participants (n=79). Among participants from this group, 36.7% preferred the American female voice (US_F), while 20.3% favored the Irish (IE_F) voice. The Indian (IN_F) voice ranked as the third most popular choice, with 17.7% of international participants preferring it. This could be attributed to the presence of Indian participants (n=52) in this group, where 26.9% preferred the Indian female voice (IN_F) and 36.7% chose the American voice (US_F). Although Ethnicity made it in the top three main themes (21.5% of responses) for female voice selections among international participants (n=79), Self-Similarity was their least relevant theme: 6.3% referred to it. Hence, user-agent voice similarity was not observed and did not emerge as a general trend among international students (n=79) when selecting female voices.
@@ 4.4. Findings: Participants’ perceptions and rationale on designers’ demographics
These findings reveal important points about the connection between the demographic characteristics of virtual humans and those of their designers. First, participants tend to relate the demographics of virtual humans to those of their designers–designer-agent demographic similarity, particularly in regard to ethnicity and gender. Considering how participants were co-designers in the first part of the study (Section 3), these findings suggest that computer science students acting as designers could be influenced by their own demographics when creating virtual humans.
Second, these findings show that the mere awareness of the designer’s demographic characteristics is insufficient for a participant to accurately identify who created the virtual human. This highlights the limitations of demographic information in determining the creators of virtual human designs.
Finally, it was observed that the visual cues of a virtual human could overshadow verbally shared information in their message, as seen with the Intervention Group - “Explicit Legacy” (n=78) who had difficulty recalling the designer’s demographics initially provided. This finding implies that the visual aspects of virtual human demographics could have a stronger impact on users’ perceptions than verbal cues, especially when agent-designer demographic similarity is present in participants’ rationale when determining the designer’s characteristics.
@@ 4.5. Findings: Designer cues and participants’ intentions towards gratitude journaling
Second, this work’s findings suggest that the demographic information of the designers, which the participants were aware of, did not provide sufficient cues to accurately identify the designer of the virtual human. Thus, this awareness may not have been strong enough to influence participants’ intentions toward promoting gratitude journaling.
Lastly, the visual cues of the virtual humans could have overshadowed the information verbally communicated by the virtual human regarding the designer’s demographics. The fact that the Intervention Group - “Explicit Legacy” (n=78) struggled to consistently remember the designer’s demographics initially provided could serve as an indication that the visual aspects of the virtual human’s demographics might have dominated their perceptions, affecting their intentions toward gratitude journaling.
Overall, these findings suggest that tailoring designer cues in virtual human design is a complex task. The visual characteristics of virtual humans and users’ expectations (or biases) could influence users’ intentions beyond the effects of any demographic similarities between users and designers; even with virtual humans explicitly communication their designer’s demographics.
@@ 5. Discussion and Future Work
This section is structured to discuss our findings in regard to the three research questions initially posed in Section 1:
  • RQ-“Similarity Elucidation”: How do similarity expectations influence the design decisions made by computer science students when co-designing virtual human for mental health support?
@@ 5.1. Insights on agent-designer similarity in virtual humans co-design
Responding to the research question RQ-“Similarity Elucidation”: How do similarity expectations influence the design decisions made by computer science students when co-designing virtual humans for mental health support?, this work’s findings provide valuable insights into the design of virtual humans to support college students in mental health conversations, especially those majoring in computer science. From the different factors identified in Section 3.4, age was found to be an essential design factor, regardless of participants’ genders or ethnic groups, for the design of virtual humans to support mental health. Although agent-student age matching was not as predominant as other demographic characteristics (e.g., place of origin), it was observed that participants’ rationale generally considered age for both male and female appearance selection (Sections 3.6.2 and 3.7.2). Hence, the appropriate tailoring of age can help promote effective virtual human interventions in assisting computer science students (and other potential populations) in mental health conversations.
@@
Regarding participants’ preferences for virtual human voices, it was observed that appearance-voice matching played a crucial role, particularly in the design of male virtual humans (Section 3.6.2). This observation was particularly observed among domestic participants, for whom ethnicity held greater significance compared to international participants. They prioritized appearance-voice matching, voice clarity, and voice comfort (Table 5). These findings support previous work by Feijóo-García et al. (Feijóo-García et al., 2023), stating the importance of appearance-voice matching in virtual humans’ effectiveness in promoting mental health conversations.
@@
This work’s findings have significant implications for designing virtual humans to support college students in sensitive contexts such as mental health. Age, gender, ethnicity, appearance-voice matching, and user-agent similarity are essential factors. Virtual human designers should aim to understand the preferences and expectations of their target population (e.g., computer science students) when creating virtual humans. Correctly tuning characteristics and agent-user similarity can determine the success of virtual human interactions.
It is important to acknowledge the limitations of this work, as the findings are derived from a specific sample of participants in one major North American college. Future research should aim to replicate and expand upon these findings with more extensive and diverse sample populations to understand better the complex interplay between age, gender, ethnicity, appearance, and voice in virtual human design for mental health contexts.
These findings contribute to understanding factors influencing computer science students’ preferences and expectations regarding virtual humans in mental health conversations. The insights gained from this work can inform the design and development of virtual humans that effectively support and engage students in meaningful mental health discussions.
@@ 5.2. Insights on the designer’s legacy and users’ perceptions of it in virtual human interactions
Responding to the research question RQ-“Designer Cues”: Which characteristics of a virtual human’s designer can be identified by computer science students during their interactions with the agent?, this work’s findings provide valuable insights into the connection between the demographic characteristics of virtual humans and those of their designers (Section 4.4). One key finding is that participants tended to associate virtual humans’ demographics with their designers’ demographics, particularly regarding ethnicity and gender. Although relevant in the selection of virtual human appearances (Sections 3.6.2 and 3.7.2), the designer’s age was not identifiable nor recognizable through the virtual human. Same for the place of origin: the characteristic that most likely presented a similarity-attraction effect (Section 3.4). Considering how participants were co-designers in the first part of the user study (Section 3), these findings suggest that computer science students acting as designers could be influenced by their own demographics when creating virtual humans.
It was observed that the mere awareness of the demographic characteristics of the virtual human’s designer was insufficient for participants to accurately identify the virtual human’s designer (Section 4.4). As reported in Section 4.4, participants’ perceptions led them to perceive the designer as demographically similar to their agent–designer-agent demographic similarity. Moreover, we observed a more intriguing phenomenon where the visual cues of the virtual human appeared to overshadow the verbally shared information regarding the designer’s demographics. Specifically, participants who interacted with a virtual human that verbally shared the demographic information of its designer had a tough time recalling the designer’s information. This finding suggests that the visual aspects of virtual human demographics hold more influence over users’ perceptions compared to verbal cues, and it may be more pronounced when participants base their rationale on agent-designer similarity. This highlights the limitations of demographic information in determining the creators of virtual human designs.
These findings contribute insights to help tailor designer cues in virtual humans for mental health support. As observed, user-designer and agent-designer similarity interact in a complex manner, even when computer science students serve as designers and users of virtual humans. Future research should explore different ways to present designer cues, adding factors apart from demographic characteristics, appearances, and voices.
@@ 5.3. Insights on the designer’s legacy effectiveness in virtual human interactions
We observed that the visual cues of the virtual humans could have overshadowed the information verbally communicated by the virtual human regarding the designer’s demographics. The fact that the Intervention Group - ”Explicit Legacy” (Section4.2.1) struggled to recall the designer’s demographics initially provided by the virtual human could serve as an indication that the visual aspects of the virtual human’s demographics might have dominated their perceptions, potentially affecting their intentions toward gratitude journaling. It was also observed that the demographic information of the designers, which the participants were aware of, did not provide sufficient cues to accurately identify the virtual human’s designer. Thus, this awareness could not have been strong enough to influence participants’ intentions toward promoting gratitude journaling.
Overall, our research suggests that tailoring designer cues in virtual human design is a complex task. This, as the visual aspects of virtual humans and users’ biases could influence users’ intentions regardless of 1) the demographic similarity between the user and the designer and 2) the virtual human explicitly communicating the designer’s demographics. Future research is needed to understand how to tailor designer cues and assess their impact on promoting intentions and behaviors toward mental health conversations.

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 42journalnumber: 1article: 1publicationmonth: 12ccs: Computing methodologies Intelligent agents
@@ 3.1. Participants
Participants’ reported ages were: 18-20 years of age (n=339), 21-25 years of age (n=119), 26-30 years of age (n=8), and over 30 years of age (n=13). Two participants (n=2) did not disclose their age. Participants also reported the number of languages they were familiar with–monolingual (n=197) or multilingual (n=284), and their place of origin–country and city. Participants reported being from 34 different countries: from the United States (n=354) or other places (n=127)–see Appendix A, Table 9. Participants also reported being domestic (n=402) or international (n=79) students.
@@ 3.3. Data analysis
Moreover, the study examined how participants selected virtual humans’ appearances and voices when prompted to choose options that would encourage ”someone similar to them” to engage in mental wellness conversations. Top preferences on appearances and voices were analyzed using descriptive statistics, cross-referenced with participants’ demographics. We used the Pearson correlation coefficient to compare participants’ self-reported ages with the perceived ages of the agents’ appearances they selected. Additionally, we calculated the Jaccard similarity coefficient to evaluate the relationship between participants’ self-reported ethnic groups and the ethnic groups they perceived the agents’ appearances they selected. Likewise, we used the Chi-Squared Test to evaluate the relationship between participants’ places of origin and their voices’ selection (Franke et al., 2012). We centered our analysis on participants who indicated being originally from India (n=57) or the United States (n=354), as the catalog of female and male voices explicitly featured American and Indian-accented voices (see Appendix A: Tables 15 and 16).
@@ 3.4. Findings: Agent-designer matching of demographic characteristics
Of the four characteristics matched between participants and virtual humans they co-designed, age was the only demographic characteristic for which agent-participant (i.e., agent-designer) matching was not the preferred choice among most participants when designing a virtual human to provide mental health assistance to someone similar to them. Three participants (n=3) did not report a counselor’s age. Among the remaining participants who self-reported their age (n=476), 66.8% of participants preferred an older virtual human (more than five years older), 0.4% preferred a younger virtual human (less than five years younger), and 32.8% designed a virtual human as old as they reported to be (having a window of ±5 years of age for similarity): virtual humans’ average age was of 28.6 years of age, reporting a maximum of 60 years of age. The Pearson correlation coefficient revealed a significant association between participants’ self-reported ages and the virtual agents’ ($r=0.257,n=476,p<0.01$), suggesting a positive relationship between them. This correlation suggests that as the participant’s age increases, the agent’s age also tends to increase, which supports the existence of user-agent age similarity when participants design agents to support individuals similar to them in mental well-being.
@@
From our sample population (n=481), 61.1% of participants decided on ethnic matching and 76.9% on matching their country of origin: 24.1% indicated the same city of origin. The sample population had male (n=281), female (n=182), and participants from other gender groups (n=18)–e.g., non-binary, transgender, etc. From the general population, 53.0% of them preferred a female agent, 37.2% a male one, and 9.8% did so for an agent with another gender identity. See Appendix A: Table 10 shows preferences on gender concerning participants’ ethnic groups.
Looking at gender similarity, 72.8% of the general population decided on gender matching. Female participants (n=182) highly preferred a female agent: 90.1% of the female population. Similarly, 83.3% of participants who identified with other gender identities (n=18–e.g., non-binary, gender-queer, transgender) preferred an agent who gender-matched them. Male participants (n=281) also leaned toward a gender-matching agent, although not as pronounced as participants from other gender groups (61.9% of males). Therefore, gender similarity is critical in the design of virtual humans to support non-male minorities in mental health.
@@ 3.6.1. Perceptions of age and ethnicity of virtual human male appearances
Of the seven ethnic groups considered, two of them did not tend toward ethnic similarity concerning male appearances. Among Latin American participants (n=100), only 33.0% of them selected a male appearance they perceived as Latin American. Likewise, only 21.4% of Middle Eastern/North African participants (n=14) leaned toward ethnically similar male appearances. As 93.0% of the Latin American and 92.3% of the Middle Eastern/North African population were domestic, further research is needed to understand if this phenomenon is due to limitations on the catalog of appearances or familiarity aspects regarding the under-representation of these ethnic groups in counseling scenarios in the United States–counselors in the United States are predominantly White (non-Latin American) (USA, 2017).
@@ 3.9. Findings: Preferences on encouraging virtual human male voices
Similar to the findings reported by Feijóo-García et al. (Feijóo-García et al., 2023), there were voice preferences among Indian participants (n=57) based on user-agent matching regarding the place of origin. There was a significant positive correlation between Indian participants’ place of origin and the selection of the female Indian voice ($\chi^{2}=79.6,df=1,p<0.01$): 63.2% of Indian participants (n=57) selected the male Indian voice (IN_M) in their top four most encouraging voices. Participants from the United States (n=354) were also observed to include the male American voice (US_M) in their four most encouraging female voices (50.3% of responses). Similarly, we found a significant correlation between American participants’ place of origin and their selection of the male American voice ($\chi^{2}=11.5,df=1,p<0.01$). However, when asked to select the most encouraging voice among each participant’s top four male voices, the male Indian voice (IN_M) was not favored among Indian participants (n=57): Only 5.3% of them preferred it as their top choice. It was the same for American participants (n=354), with only 21.5% of them selecting the male American voice (US_M) as the most preferred voice among the top four male voices each participant selected.
@@ 3.10. Findings: Preferences on encouraging virtual human female voices
Similar to observations on participants’ choices of male voices (Section 3.9), Indian participants (n=57) leaned toward female voices that matched their place of origin (i.e., India) when selecting the four most encouraging female voices. There was a significant positive correlation between Indian participants’ place of origin and the selection of the female Indian voice ($\chi^{2}=94.5,df=1,p<0.01$): 80.7% of Indians (n=57) selected the female Indian voice (IN_F) in their top four most encouraging voices. Participants from the United States (n=354) were also observed to include the female American voice (US_F) in their four most encouraging female voices (56.2% of responses). However, contrary to what we observed among Indian participants, there was not a significant correlation between American participants’ place of origin and their selection of the female American voice ($\chi^{2}=3.22,df=1,p=0.07$). On a different direction, when asked to select the most encouraging voice among the top four female voices each participant selected, Indian participants (n=57) ended up selecting voices different from the female Indian voice (IN_F): Only 29.8% of them preferred it. It was the same for American participants (n=354), with only 24.9% of them selecting the female American voice (US_F) as the most preferred voice among the top four female voices each participant selected.
@@
Accent similarity expectations were not observed among international participants (n=79) when selecting their top-most encouraging female voices. Among participants from this group, 36.7% preferred the American female voice (US_F), while 20.3% favored the Irish (IE_F) voice. The Indian (IN_F) voice ranked as the third most popular choice, with 17.7% of international participants preferring it. This could be attributed to the presence of Indian participants (n=52) in this group, where 26.9% ranked first the Indian female voice (IN_F) and 36.7% preferred the American voice (US_F). Although Ethnicity made it in the top three main themes (21.5% of responses) for female voice selections among international participants (n=79), Self-Similarity was their least relevant theme: 6.3% referred to it. Hence, user-agent voice similarity was not observed and did not emerge as a general trend among international students (n=79) when selecting female voices.
@@ 5. Discussion and Future Work
This section is structured to discuss our findings in regard to the three research questions initially posed in Section 1.
@@ 5.1. Insights on agent-designer similarity in virtual humans co-design
Responding to the research question RQ-“Similarity Elucidation”: How do similarity expectations influence the design decisions made by computer science students when co-designing virtual humans for mental health support?, this work’s findings provide valuable insights into the design of virtual humans to support college students in mental health conversations, especially those majoring in computer science. From the different factors identified in Section 3.4, age was found to be an essential design factor, regardless of participants’ genders or ethnic groups, for the design of virtual humans to support mental health. Although user-agent age matching was not as predominant as other demographic characteristics (e.g., place of origin), it was observed that participants’ rationale generally considered age for both male and female appearance selection (Sections 3.6.2 and 3.7.2), suggesting age can help promote effective virtual human interventions in assisting computer science students (and other potential populations) in mental health conversations.
@@
Regarding participants’ preferences for virtual human voices, it was observed that appearance-voice matching played a crucial role, particularly in the design of male virtual humans (Section 3.6.2). This observation was particularly seen among domestic participants, for whom ethnicity had more prevalence compared to international participants. They prioritized appearance-voice matching, voice clarity, and voice comfort (Table 5). These findings support previous work by Feijóo-García et al. (Feijóo-García et al., 2023), stating the importance of appearance-voice matching in virtual humans’ effectiveness in promoting mental health conversations.
@@
These findings contribute to understanding factors influencing computer science students’ preferences and expectations regarding virtual humans in mental health conversations. Age, gender, ethnicity, appearance-voice matching, and user-agent similarity were found to be essential factors. Virtual human designers should aim to understand the preferences and expectations of their target population (e.g., computer science students) when creating virtual humans. The insights gained from this work can inform the design and development of virtual humans that effectively support and engage students in meaningful mental health conversations. Future research should aim to replicate and expand upon these findings with more extensive and diverse sample populations to understand better the complex interplay between age, gender, ethnicity, appearance, and voice in virtual human design for mental health contexts.
@@ 5.2. Insights on the designer’s legacy and users’ perceptions of it in virtual human interactions
Responding to the research question RQ-“Designer Cues”: Which characteristics of a virtual human’s designer can be identified by computer science students during their interactions with the agent?, this work’s findings provide valuable insights into the connection between the demographic characteristics of virtual humans and those of their designers (Section 4.4).
First, participants tend to relate the demographics of virtual humans to those of their designers–designer-agent demographic similarity, particularly in regard to ethnicity and gender. Although relevant in the selection of virtual human appearances (Sections 3.6.2 and 3.7.2), the designer’s age was not identifiable nor recognizable through the virtual human. Same for the place of origin: the characteristic that most likely presented a similarity-attraction effect (Section 3.4). Considering how participants were co-designers in the first part of the user study (Section 3), these findings suggest that computer science students acting as designers could be influenced by their own demographics when creating virtual humans.
Second, we observed that the mere awareness of the demographic characteristics of the virtual human’s designer was insufficient for participants to accurately identify the virtual human’s designer (Section 4.4). As reported in Section 4.4, participants’ perceptions led them to perceive the designer as demographically similar to their agent–designer-agent demographic similarity.
Finally, we observed a more intriguing phenomenon where the visual cues of the virtual human appeared to overshadow the verbally shared information regarding the designer’s demographics. Specifically, participants who interacted with a virtual human that verbally shared the demographic information of its designer (see Section 4.2.1: Intervention Group - “Explicit Legacy”) had a tough time recalling the designer’s information. This observation suggests that the visual aspects of virtual human demographics hold more influence over users’ perceptions compared to verbal cues, and it may be more pronounced when participants base their rationale on agent-designer similarity. This highlights the limitations of demographic information in determining the creators of virtual human designs. User-designer and agent-designer similarity interact in a complex manner, even when computer science students serve as designers and users of virtual humans. Future research should explore different ways to present designer cues, adding factors apart from demographic characteristics, appearances, and voices.
@@ 5.3. Insights on the designer’s legacy effectiveness in virtual human interactions
Second, this work’s findings suggest that the demographic information of the designers, which the participants were aware of, did not provide sufficient cues to accurately identify the designer of the virtual human. Thus, this awareness may not have been strong enough to influence participants’ intentions toward promoting gratitude journaling. Moreover, the fact that the Intervention Group - “Explicit Legacy” (Section4.2.1) struggled to consistently remember the designer’s demographics initially provided could serve as an indication that the visual aspects of the virtual human’s demographics might have dominated participants’ perceptions, potentially affecting their intentions toward gratitude journaling. The visual aspects of virtual humans and users’ biases could influence users’ intentions regardless of 1) the demographic similarity between the user and the designer and 2) the virtual human explicitly communicating the designer’s demographics. Future research is needed to understand how to tailor designer cues and assess their impact on promoting intentions and behaviors toward mental health conversations.